OpenStudy (anonymous):

Part a.) Write the equation of the line that has a slope of -2 and a y-intercept of -6. (5 points) Part b.) Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3. (5 points)

OpenStudy (lalaly):

y=mx+c where m is the slope and c is the y intercept

OpenStudy (anonymous):

a) y=kx+m where k=slope and, m=y-intercept --> y=-2x-6

OpenStudy (anonymous):

b) same principle... y=kx+m --> -2=-2/3*6+m m=2 --- > y=-2/3x+2 Hope this helps... /Ali